OpenStudy (anonymous):

PROBLEM FROM HELL! HOW CAN I FIND X? OH JESUS! \[\Huge \dfrac{ x^{\pi} + x^x - x^{4x^2-3x+123} } {3^x + x^{3^{5x}}} = 0\]

OpenStudy (anonymous):

ITS flutterING TOO HARD!

OpenStudy (anonymous):

SOMEBODY HELP ME PLEASE!

OpenStudy (anonymous):

@yummydum help me please.... im dying here..... oh my god.....

OpenStudy (anonymous):

x^pi + x^x + x^(4x^2-3x+123) = 0

OpenStudy (anonymous):

i dont know how to do this sorry....try plugging it into http://www.wolframalpha.com/ ...that should help probably :)

OpenStudy (anonymous):

Hint: For A/B to be 0, A must be 0 and B cannot be 0.

OpenStudy (anonymous):

HOW THE [BLEEP] CAN I SOLVE x^pi + x^x + x^(4x^2-3x+123) = 0!?!?!?

OpenStudy (anonymous):

idk, most of time w exponents you take ln of both side, but you cant take ln of 0 in this case...

OpenStudy (anonymous):

nvm i think i got it

OpenStudy (anonymous):

@Sheng This question is closed now. ;)

OpenStudy (anonymous):

x^pi + x^x = x^(4x^2-3x+123) ln both side pilnx + xlnx = (4x^2-3x+123)lnx

OpenStudy (anonymous):

aweh ok i was having fun w it too D:

OpenStudy (anonymous):

Ah, great! I don't think ... I don't *THINK* it is possible to do this. Good job, problem writer.